A plane of two turns coil of area 0.0127 m2 is placed in a uniform magnetic field of 0.7 T. The current in the coils is 3 A. If the plane makes an angle of 50° with the magnetic field, calculate the torque exerted on the coil.


1
Expert's answer
2022-02-23T12:11:23-0500

Answer

torque exerted on the coil

"\tau=NIABsin\theta"

Putting all values

"\tau=2*3*0.0127*0.7*\sin50°\\=253.34Nm"